OpenStudy (anonymous):

LHospital Rule lim x->infinity (1=(2/x^2))^lnx

OpenStudy (anonymous):

I can't make sense of this

OpenStudy (anonymous):

nor me!

OpenStudy (anonymous):

\[\lim x-> \infty (1+2/x ^{2})^{lnx}\]

OpenStudy (anonymous):

as x --> inf 2 /x^ .---> 0 and 1 to any power = 1 limit is 1

OpenStudy (zarkon):

1 is the answer...but your method doesn't always work

OpenStudy (alfie):

Using L'hopital you should first get to one of the forms allowed from the rule. What you can do is apply a function and its inverse so that everything remains unchanged but it gives you the opportunity to go to a form such as infinity/infinity. I usually use the exponential and the logarithm and then use them properties.

OpenStudy (alfie):

woops... their* . English mistake :P

OpenStudy (anonymous):

please explain zarkon

OpenStudy (zarkon):

@ jimmyrep compute this limit \[\lim_{x\to\infty}\left(1+\frac{2}{x^2}\right)^{x^2}\]

OpenStudy (zarkon):

I'll give you a hint...the answer is not 1

OpenStudy (anonymous):

you've got me there - limits are'nt my strong point!

OpenStudy (jamesj):

For the original problem--but not Zarkon's, which is very important-- consider the log of the expression: \[ \ln (1+2/x^2)^{\ln x} = \ln x . \ln(1 + 2/x^2) = \frac{\ln(1 + 2/x^2)}{ 1/\ln x} \]

OpenStudy (jamesj):

Now you can apply l'Hopital's rule.

OpenStudy (anonymous):

aahh zarkons problem its the exponential! limit is e^x right?

OpenStudy (jamesj):

Yes, \[ \lim_{x \rightarrow \infty} (1 + a/x)^{1/x} = e^a \]

OpenStudy (anonymous):

thanks for confirming this - i a[ways have trouble with limits. to me they are most difficult part of calculus